OpenStudy (anonymous):

I need to do an explanation of a graph family and I dont understand what is being asked. Graph y=a(absolute value of x)+b.

OpenStudy (phi):

Did they give you numbers for a and b ?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

I am supposed to draw it and explain what happens, what I notice, what changes..

OpenStudy (phi):

a and b can be anything. you could start with a=0 b=0 and a= 1 b=0 you get a "V" a= -1 b=0 upside down V then try with b= 1 and then b= -1 (the V's move up or down) then try with a=1/2 and then a= 2 that should be enough

OpenStudy (anonymous):

Thanks, will give that a try, just an odd question that is not covered at all in chapter.